# YITP-24-178

### v1

@@ 1 Introduction
The standard $\Lambda$CDM model of cosmology provides a comprehensive explanation for cosmological and astrophysical observations. According to this model, the matter content of the Universe consists of the cosmological constant, dark matter, and ordinary matter that includes baryonic matter and radiation. The cosmological constant $\Lambda$ is more generically called dark energy by allowing its time variation. The existence of dark energy is necessary to explain the observed late time cosmic acceleration [1, 2]. The current understanding is that dark energy constitutes about 70% of the energy content of the Universe. The dark matter, which is needed to explain the observed rotation curves of the galaxies as well as the structure formation in the Universe [3, 4, 5], contributes about 27%, while the visible sector, consisting of the ordinary baryonic matter and radiation, contributes only 3% of the energy content of the Universe [6]. Although the $\Lambda$CDM model is observationally successful, it has its own shortcomings from the theoretical point of view. We have no compelling theoretical explanation for the current value of the cosmological constant $\Lambda$ or the dark energy density, which is called the cosmological constant problem [7, 8]. A more subtle mystery is the origin of dark matter, which remains unknown.
@@
As dark energy and dark matter constitute around 97% of the content of the Universe, it is very tempting to consider a scenario in which the dark energy and/or dark matter are also produced during inflation. However, this is not an easy task. Thus, most of scenarios deal with the production of the dark energy and dark matter particles after inflation, i.e. during the radiation dominance. Nevertheless, there has been some efforts to construct scenarios for the production of dark energy [11, 12, 13, 14] and dark matter particles [15, 16, 17, 18] during inflation.
@@ 2 The model
We are interested in quantum fluctuations of the field $X$, assuming that the vacuum expectation value (vev) of $X$ is negligible. In cosmological background, the modes with wavelengths larger than the Hubble horizon can be treated as a homogeneous vev and one naturally expects $\langle{{\hat{X}}}\rangle\neq 0$ in our observable universe.111 Here we have placed ${}{\hat{}}~{}$ to make it clear that the field is an operator. The state for which the expectation values are computed is assumed to be the one annihilated by the annihilation operators. Thus the proper statement of our assumption will be
@@
we expand the Fourier amplitude as ${\hat{X}}_{\bf k}(\tau)=X_{k}(\tau)\,\hat{a}_{\bf k}+X^{\ast}_{k}(\tau)\,\hat{a}_{-\bf k}^{\dagger}$ where $X_{k}(\tau)$ is the mode function, $\hat{{a}}_{\bf k}$ and $\hat{{a}}_{\bf k}^{\dagger}$ are annihilation and creation operators which satisfy the usual commutation relations $[\hat{{a}}_{\bf k},\hat{{a}}_{\bf q}^{\dagger}]=\delta({\bf k}-{\bf q})$. We then find the mode function satisfies
@@
$\Omega_{X}(k,\tau)\equiv\frac{1}{3M_{\rm Pl}^{2}H^{2}}\frac{{\rm d}{\bar{\rho}}_{X}}{{\rm d}\ln{k}}=\frac{f^{2}k^{3}}{12\pi^{2}M_{\rm Pl}^{2}a^{2}H^{2}}\left[\big{|}X^{\prime}_{k}\big{|}^{2}+\left(c_{s}^{2}k^{2}+m^{2}a^{2}\right)\big{|}X_{k}\big{|}^{2}\right]\,.$ (2.9)
@@ 3 Superhorizon modes during inflation
In our scenario, the initial condition is set during inflation. In order to make it possible for the quantum fluctuations of $X$ to be frozen on superhorizon scales, we need to assume that $X$ is a light field during inflation,
@@
[…]
@@
for some period during inflation. This condition gives an upper limit $k_{\max}$ on $k$, and any modes with $k\lesssim k_{\rm max}$ acquire perturbations which eventually freeze. The explicit value of $k_{\max}$ can be determined for a given functional form of $f$. For instance, if $f=1$, the condition (3.3) is satisfied only for the superhorizon modes $c_{s}{k}\ll{a}H_{\rm inf}$ or $-c_{s}{k}\tau\ll 1$ where we have used $a\simeq-1/(H_{\rm inf}\tau)$. Here we mention that by horizon we mean the sound horizon $c_{s}k=a^{\prime}/a$. If $f=a^{-1}$, which happens for spin-1 fields, then it is not possible to satisfy (3.3). This is a well-known fact that spin-1 fields will not feel the de Sitter horizon due to the conformal symmetry. In that case, having $f\neq a^{-1}$ (non-minimal coupling which can be achieved, i.e., through direct interactions with the inflaton field) is necessary to break the conformal symmetry and generate an effective horizon for vector modes [52, 33, 37]. Thus, for $f\neq a^{-1}$, condition (3.3) can be satisfied for the modes which are not necessarily superhorizon.
@@
[…]
@@ 4.1 Relativistic regime
and $p_{s}(z)$ and $q_{s}(z)$ are defined in (A.2) and (A.3). Note that the series form (4.10) is valid for any values of $\mu$ as long as $\mu_{k}\ll 1$. Also note that ${\rm arg}[a]\to\pi/2-\epsilon$ and ${\rm arg}[z]=\pi/2$. Thus, we have $2ia^{\frac{1}{2}}z^{\frac{1}{2}}=-x\left(1-3i\mu_{k}\right)^{1/2}$, hence the argument of trigonometric functions in (4.10) becomes real and equal to $-x=-c_{s}k$ in the limit $\mu_{k}\ll 1$.
@@ 5.1.1 $m\ll{H}$ ($\Leftrightarrow\mu\ll 1$)
$\begin{split}\Omega_{X}(k,\tau)&=\frac{{\cal P}_{X,i}(k)}{6M_{\rm Pl}^{2}}\begin{cases}x^{2}\left(1+\mu_{k}^{2}\,x^{2}\right)\,;&x<1\,,\\ 1-\dfrac{\sin(2x)}{x}\left(1+\dfrac{1}{3}\mu_{k}^{2}\,x^{2}\right)+\dfrac{1}{6}\mu_{k}^{2}\,x^{2}\cos(2x)\,;&x\gg 1\,,\end{cases}\end{split}$ (5.2)
@@ 5.2.1 $m\ll{H}$ ($\Leftrightarrow\mu\ll 1$)
up to a small correction proportional to the inverse of mass squared. This means the energy density ${\bar{\rho}}_{X}(\tau)$ of these superhorizon modes is constant in time.
@@ 5.2.2 $m\gg{H}$ ($\Leftrightarrow\mu\gg 1$)
which shows that the energy behaves like dust. Note that this result holds for both subhorizon and superhorizon modes.
@@ 6.1 Amplitude of the dark matter density contrast
From now on, we set the sound speed to be unity $c_{s}=1$ for the sake of simplicity. As the CMB and LSS scales are more or less around $k_{\rm eq}$, where $k_{\rm eq}\sim{10}^{-2}\mbox{Mpc}^{-1}$ is the comoving wavenumber which re-enters the horizon at the time of matter-radiation equality, we should restrict our scenario to the case where the scale-dependent initial power spectrum with $f\neq 1$ excludes excitations of the modes $X_{k}$ with $k\lesssim{k}_{\rm eq}$. Actually, in this case, the dimensionless power spectrum of $\delta_{X}$ is extremely blue on large scales. It scales as ${\cal P}_{\delta_{X}}\propto k^{3}$ in general, except for the case of a monochromatic power spectrum proportional to Dirac’s delta function in which $\propto k^{2}$. Thus, adiabatic perturbations coming from the inflaton fluctuations, which are nearly scale-invariant, become a dominant component of dark matter perturbations on large scales and there is no deviation from the $\Lambda$CDM model predictions on those scales.666 Adiabatic perturbations from the inflaton are imprinted in the perturbations of $X$ at the onset of its oscillations. This is because the starting time of the oscillations is determined by the uniform density hypersurface ($H=m$) on which the perturbation of radiation density vanishes.
@@
where ${\cal A}$ is a free dimensionless coefficient. Thus, in this simplest setup, there are three free parameters in the model; the mass $m$, the scale $k_{p}$, and the amplitude ${\cal A}$.
@@ 6.2 Relic density
in which $r={\cal P}_{h}/{\cal P}_{\cal R}$ is the tensor-to-scalar ratio, ${\cal P}_{h}=2H_{\rm inf}^{2}/(\pi^{2}M_{\rm Pl}^{2})$ and ${\cal P}_{\cal R}$ are the power spectra of the primordial gravitational waves and the curvature perturbation, respectively. Note that since $X$ is a spectator field during inflation, it does not significantly contribute to the metric tensor perturbation and curvature perturbation. Thus, ${\cal P}_{h}$ and ${\cal P}_{\cal R}$ are the same as the usual power spectra in single field inflation. However, in the case that $X$ is a spin-2 field, there can be a significant contribution to ${\cal P}_{h}$ [55, 56]. In this case, one needs to carefully take into account the contribution of the field $X$ as a source of gravitational waves.
@@ 6.3 Constraints on the model parameters
On the other hand, finding a lower bound on the mass needs careful considerations since it completely depends on the scale of our interest. However, as we will show, under the assumption of a sharply peaked power spectrum at $k_{p}$, it can be done in a simple manner. In what follows, in addition to the above constraint on the mass, we give four conditions based on the cosmological observations that support the standard $\Lambda$CDM model. We will use them to identify the allowed region in the parameter space.
@@ 6.4 Halo formation
Although any point in the allowed region is equally allowed at the level of the conditions we have imposed, our model may be tested against future observations, since it predicts characteristic inhomogeneities on small scales that do not exist in the standard $\Lambda$CDM model. One notable feature is that the dark matter density contrast at the scale $k_{p}$ has its variance $\langle\delta_{X}^{2}\rangle=2$ from the outset. This magnitude is much larger than the one $\sim 10^{-10}$ observed at CMB and LSS scales. Such large dark matter fluctuations will result in the copious formation of dark matter halos in the early times even around the time of the matter-radiation equality.
In order to qualitatively estimate in which region in the parameter space this can happen, let us compare the two time scales, the sound crossing time $t_{s}$ and the free-fall time $t_{f}$, of an overdense region with its comoving size $1/k_{p}$. Because the proper wavenumber is $k_{p}/a$, the propagation speed of the wave is $\frac{k_{p}}{am}$. Thus, the sound crossing time, which is the time the waves take to propagate over the size $a/k_{p}$, becomes $t_{s}=\frac{a^{2}m}{k_{p}^{2}}$. The free-fall time is $t_{f}=1/\sqrt{G\rho}$ ($\sim M_{\rm Pl}/\sqrt{\rho}$ in our unit). Gravitational collapse of the overdense region will occur when $t_{f}<t_{s}$ is satisfied. Thus, the modes satisfying
@@
After the time of matter-radiation equality, the redshift range is $0\leq{z}\lesssim 3000$ which gives $5.50.1\leq{\left(\frac{1+z}{100}\right)}^{\frac{1}{2}}<$ 5.5. Thus, the condition to form dark matter halo is approximately defined by $m>H_{p}$ which corresponds to the modes which are non-relativistic at the time of horizon re-entry.
@@
Thus, only subsolar mass dark matter halos can form from the large isocurvature perturbations of $X$. Applying the CMB and LSS bound (6.16) to (6.21) and using the lower bound on the mass (6.17), we find that within the mass range $10^{-21}\,\mbox{eV}\lesssim m<10^{-16}\,\mbox{eV}$ dark matter halos cannot form. For the mass range $m\gtrsim 10^{-16}\,{\rm eV}$ subsolar mass dark matter halos will form at high redshifts (see Fig. 3). Observational confirmation of a large population of such low-mass dark matter halos will support our scenario that dark matter originated purely from the quantum fluctuations in the inflationary period.
@@ 7 Summary and discussion
We studied a scenario in which dark matter is a massive bosonic field, arising solely from the quantum fluctuations that are stretched to the superhorizon scales during inflation. In this scenario, dark matter exhibits ${\cal O}(1)$ primordial isocurvature perturbations at small scales that are beyond the reach of current observations, such as those from the CMB and LSS. Neglecting changes in the number of relativistic degrees of freedom, we derived an exact transfer function for the dark matter field perturbations during the radiation dominated era. Depending on the hierarchy between the mass $m$, the scale $k$, and the Hubble parameter $H$, the perturbations show distinct behaviors across different regimes. We derived approximate expressions for the transfer function for each regime and reproduced the known behaviors discussed in the literature. Assuming a monochromatic initial power spectrum with a peak at $k=k_{p}$, we identified the viable parameter space in the $m$-$k$ plance as $10^{-21}\,\mbox{eV}\lesssim{m}\ll H_{\rm inf}$ and $k_{p}\gtrsim 10~{}{\rm kpc}^{-1}$. The lower bound on the mass comes from the observation of the galactic structures while the upper bound should be imposed to have effectively massless perturbations during inflation such that they freeze at the superhorizon scales. The lower bound on the scale $k_{p}$ is to avoid large fluctuations, e.g. large isocurvature perturbations, on large scales probed already by the CMB and LSS observations.
@@
Finally, let us mention possible mechanisms to realize our scenario. The most important assumption of our scenario is that the field $X$ is light during inflation, i.e., $m\ll H_{\rm inf}$. Such a light field during inflation may be realized in models based on the symmetry breaking mechanism: the field can be massless during inflation and it acquires mass at the end of inflation or even after inflation through the symmetry breaking mechanism [36, 37]. However, achieving the nearly massless condition for a spectator field with negligible vev (2.2) may not be easy as the stochastic motion generally leads to non-zero $\langle\hat{X}\rangle$ in our observable universe. The common way to achieve (2.2) is to have a potential that makes $X$ settle down to $X=0$, but it requires a large mass which contradicts our assumption. One of the possibilities is to consider a two-stage inflationary scenario in which $X$ is pinned to $X=0$ through the large mass at the first stage and becomes almost massless at the second stage. Whether such a model can be constructed in a natural manner is a question to be answered in future work.

### v2

@@ 1 Introduction
[…]
@@
As dark energy and dark matter constitute around 97% of the content of the Universe, it is very tempting to consider a scenario in which the dark energy and/or dark matter are also produced during inflation. However, this is not an easy task. Thus, most of scenarios deal with the production of the dark energy and dark matter particles after inflation, i.e. during the radiation dominance. Nevertheless, there has been some efforts to construct scenarios for the production of dark energy [12, 13, 14, 15, 16, 17] and dark matter particles [18, 19, 20, 21] during inflation.
@@ 2 The model
We are interested in quantum fluctuations of the field $X$, assuming that the vacuum expectation value (vev) of $X$ is negligible. In cosmological background, the modes with wavelengths larger than the Hubble horizon can be treated as a homogeneous vev and one naturally expects $\langle{{\hat{X}}}\rangle\neq 0$ in our observable universe. [^1] Here we have placed $\hat{}$}~ to make it clear that the field is an operator. The state for which the expectation values are computed is assumed to be the one annihilated by the annihilation operators. Thus the proper statement of our assumption will be
@@
we expand the Fourier amplitude as ${\hat{X}}_{\bf k}(\tau)=X_{k}(\tau)\,\hat{a}_{\bf k}+X^{\ast}_{k}(\tau)\,\hat{a}_{-\bf k}^{\dagger}$ where $X_{k}(\tau)$ is the mode function, $\hat{{a}}_{\bf k}$ and $\hat{{a}}_{\bf k}^{\dagger}$ are annihilation and creation operators which satisfy the usual commutation relations $\hat{{a}}_{\bf k},\hat{{a}}_{\bf q}^{\dagger}]=\delta({\bf k}-{\bf q})$. We then find the mode function satisfies
@@
[…]
@@ 3 Superhorizon modes during inflation
In our scenario, the initial condition is set during inflation. In order to make it possible for the quantum fluctuations of $X$ to be frozen on superhorizon scales, we need to assume that $X$ is a light field during inflation, [^2] One may notice that assuming (3.1) makes it difficult to achieve (2.2). This is a general feature of models based on dark matter production from frozen superhorizon perturbations. Therefore, a mechanism is required to reconcile these conditions. We propose possible scenarios in Sec. 7.
@@
In flat spacetime, where both $a$ and $f$ are constant, the second friction-like term in the l.h.s. disappears and the above equation has oscillating solution of the form $e^{\pm{i}c_{s}{k}\tau}$ which characterize the usual vacuum fluctuations. On the other hand, in the time-dependent FLRW background, the friction term can make the modes freeze. In the long-wavelength limit $c_{s}k\to 0$, two independent solutions are $X_{k}={\rm const.}$ and the time-dependent part of $X_{k}=\int^{\tau}d\tau^{\prime}/(a^{2}f^{2})$ with the latter being a quickly decaying function.
@@
for some period during inflation. This condition gives an upper limit $k_{\max}$ on $k$, and any modes with $k\lesssim k_{\rm max}$ acquire perturbations which eventually freeze. The explicit value of $k_{\max}$ can be determined for a given functional form of $f$. For instance, if $f=1$, the condition (3.3) is satisfied only for the superhorizon modes $c_{s}{k}\ll{a}H_{\rm inf}$ or $c_{s}{k}\tau\ll 1$ where we have used $a\simeq-1/(H_{\rm inf}\tau)$. Here we mention that by horizon we mean the sound horizon $c_{s}k=a^{\prime}/a$. If $f=a^{-1}$, which happens for spin-1 fields, then it is not possible to satisfy (3.3). This is a well-known fact that spin-1 fields will not feel the de Sitter horizon due to the conformal symmetry. In that case, having $f\neq a^{-1}$ (non-minimal coupling which can be achieved, i.e., through direct interactions with the inflaton field) is necessary to break the conformal symmetry and generate an effective horizon for vector modes [55, 36, 40]. Thus, for $f\neq a^{-1}$, condition (3.3) can be satisfied for the modes which are not necessarily superhorizon.
@@
where $X_{k,i}\equiv{X}_{k}(\tau_{i})$ denotes the value of $X_{k}(\tau)$ at superhorizon regime $c_{s}k\tau_{i}\to 0$. [^4] To be more precise, if we match the scale factor at the end of inflation to the radiation stage at $\tau=\tau_{i}(<0)$, the conformal time for the radiation-dominated universe will be slightly shifted as $a\propto(\tau-2\tau_{i})$. Nevertheless, for the modes of our interest $c_{s}k\tau_{i}\ll 1$, we may ignore this shift and assume $a\propto\tau$. In the rest of this paper, we call (3.4) initial power spectrum. The whole role of the coupling function $f$ is to have different shapes and scale-dependency for the initial power spectrum imprinted during inflation.
@@ 4.1 Relativistic regime
and $p_{s}(z)$ and $q_{s}(z)$ are defined in (A.2) and (A.3). Note that the series form (4.10) is valid for any values of $\mu$ as long as $\mu_{k}\ll 1$. Also note that ${\rm arg}[a]\to\pi/2-\epsilon$ and ${\rm arg}[z]=\pi/2$. Thus, we have $2ia^{\frac{1}{2}}z^{\frac{1}{2}}=-x\left(1-3i\mu_{k}\right)^{1/2}$, hence the argument of trigonometric functions in (4.10) becomes real and equal to $x=-c_{s}k$ in the limit $\mu_{k}\ll 1$.
@@ 5.1.1 $m\ll{H}$ ($\Leftrightarrow\mu\ll 1$)
[…]
@@
For $x<1$, the gradient term dominates, and $\bar{\Omega}_{X}(\tau)$ behaves like spatial curvature. In contrast, for $x\gg 1$, both kinetic and gradient terms contribute equally, and $\bar{\Omega}_{X}(\tau)$ behaves like radiation, as expected. In the former case, there may be interesting features, which are beyond the scope of this paper. In the latter case, the additional dark radiation would contribute to the effective number of relativistic degrees of freedom [7], leading to constraints from big bang nucleosynthesis (BBN).
@@ 5.1.2 $m\gg{H}$ ($\Leftrightarrow\mu\gg 1$)
Thus, in this case, we find
[…]
which shows that the averaged energy density behaves like radiation. This is because, for modes deep inside the horizon ($x\gg 1$), both kinetic and gradient terms contribute equally. Similar to the $x\gg 1$ case in (5.3), there will be a BBN constraint on the energy density.
@@ 5.2.1 $m\ll{H}$ ($\Leftrightarrow\mu\ll 1$)
up to a small correction proportional to the inverse of mass squared. This means the energy density ${\bar{\rho}}_{X}(\tau)$ of these superhorizon modes is constant in time. This is because the mass term dominates, and therefore the energy density behaves like a cosmological constant.
@@ 5.2.2 $m\gg{H}$ ($\Leftrightarrow\mu\gg 1$)
which shows that the energy behaves like dust. This is because the kinetic and mass terms dominate equally. Note that this result holds for both subhorizon and superhorizon modes. This is indeed the case of interest in this paper, as $X$ can play the role of dark matter, which is the subject of the next section.
@@ 6.1 Amplitude of the dark matter density contrast
[…]
@@
where $\cal A$} is a free dimensionless coefficient. Note that $\cal A$} can be very large such that the power spectrum is only enhanced at $k=k_{p}$ and completely suppressed at other scales [56]. Thus, in this simplest setup, there are three free parameters in the model; the mass $m$, the scale $k_{p}$, and the amplitude $\cal A$}.
@@ 6.2 Relic density
[…]
@@ 6.3 Constraints on the model parameters
[…]
@@ 6.4 Halo formation
[…]
In order to qualitatively estimate in which region in the parameter space this can happen, let us compare the two time scales, the sound crossing time $t_{s}$ and the free-fall time $t_{f}$, of an overdense region with its comoving size $1/k_{p}$. Because the proper wavenumber is $k_{p}/a$, the propagation speed of the wave is $\frac{k_{p}}{am}$. Thus, the sound crossing time, which is the time the waves take to propagate over the size $a/k_{p}$, becomes $t_{s}=\frac{a^{2}m}{k_{p}^{2}}$. The free-fall time is $t_{f}=1/\sqrt{G\rho}$ ($\sim M_{\rm Pl}/\sqrt{\rho}$ in our unit). Gravitational collapse of the overdense region will occur when $t_{f}<t_{s}$ is satisfied [65, 66]. Thus, the modes satisfying
@@
After the time of matter-radiation equality, the redshift range is $0\leq{z}\lesssim 3000$ which gives $0.1\leq{\left(\frac{1+z}{100}\right)}^{\frac{1}{2}}<5.5$. Thus, the condition to form dark matter halo is approximately defined by $m>H_{p}$ which corresponds to the modes which are non-relativistic at the time of horizon re-entry.
@@
Thus, only subsolar-mass dark matter halos can form from the large isocurvature perturbations of $X$. Applying the CMB and LSS bound (6.16) to (6.21), and using the lower bound on the mass (6.17), we find that within the mass range $10^{-21}\,\mathrm{eV}\lesssim m<10^{-16}\,\mathrm{eV}$, dark matter halos cannot form. For the mass range
$m\gtrsim 10^{-16}\,\mathrm{eV}\,,\qquad\mbox{subsolar-mass halos form}\,$, (6.25)
subsolar-mass dark matter halos will form at high redshifts (see Fig. 3). Observational confirmation of a large population of such low-mass dark matter halos would support our scenario in which dark matter originates purely from quantum fluctuations during inflation. Here we mention that the de Broglie length constraint on the halo formation could further limit the mass range of our interest [6]. However, a straightforward estimate tells us that the constraint (6.25) we obtained, is already strong enough to satisfy the de Broglie length constraint.
@@ 7 Summary and discussion
We studied a scenario in which dark matter is a massive bosonic field, arising solely from the quantum fluctuations that are stretched to the superhorizon scales during inflation. In this scenario, dark matter exhibits ${\cal O}(1)$ primordial isocurvature perturbations at small scales that are beyond the reach of current observations, such as those from the CMB and LSS. Neglecting changes in the number of relativistic degrees of freedom, we derived an exact transfer function for the dark matter field perturbations during the radiation dominated era. Depending on the hierarchy between the mass $m$, the scale $k$, and the Hubble parameter $H$, the perturbations show distinct behaviors across different regimes. We derived approximate expressions for the transfer function for each regime and reproduced the known behaviors discussed in the literature. Assuming a monochromatic initial power spectrum with a peak at $k=k_{p}$, we identified the viable parameter space in the $m$-$k$ plane as $10^{-21}\,\mbox{eV}\lesssim{m}\ll H_{\rm inf}$ and $k_{p}\gtrsim 10~{\rm kpc}^{-1}$. The lower bound on the mass comes from the observation of the galactic structures while the upper bound should be imposed to have effectively massless perturbations during inflation such that they freeze at the superhorizon scales. The lower bound on the scale $k_{p}$ is to avoid large fluctuations, e.g. large isocurvature perturbations, on large scales probed already by the CMB and LSS observations.
@@
Finally, let us mention possible mechanisms to realize our scenario. The most important assumption of our scenario is that the field $X$ is light during inflation, i.e., $m\ll H_{\rm inf}$. Such a light field during inflation may be realized in models based on the symmetry breaking mechanism: the field can be massless during inflation and it acquires mass at the end of inflation or even after inflation through the symmetry breaking mechanism [39, 40]. However, achieving the nearly massless condition for a spectator field with negligible vev (2.2) may not be easy as the stochastic motion generally leads to non-zero $\langle\hat{X}\rangle$ in our observable universe. The common way to achieve (2.2) is to have a potential that makes $X$ settle down to $X=0$, but it requires a large mass which contradicts our assumption. A simple way to overcome this difficulty is to introduce a time-dependent non-minimal coupling of the form $g(t)RX^{2}$, where $g$ depends on time through the time-dependence of the inflaton, i.e., $g(\phi)=g(\phi(t))$. This time dependence makes the effective mass $m_{\rm eff}^{2}=m^{2}+g(t)H^{2}$ larger than $H^{2}$ at the early stage when $g={\cal O}(1)$, while $g$ becomes very small $g\ll 1$ at the later stage when the scale $k_{p}$ crosses the horizon. Another possibility is to consider a two-stage inflationary scenario in which $X$ is pinned to $X=0$ due to the large mass in the first stage and becomes almost massless at the second stage. We leave explicit construction of such models, as well as the study of the effects of interactions, for future work.